OpenStudy (anonymous):

Factor the following polynomial completely. 27^3+8

OpenStudy (solomonzelman):

well, you know that 8=2³

OpenStudy (anonymous):

Yes I do.

OpenStudy (solomonzelman):

So use the rule, a³+b³=(a+b)(a²-ab+b²)

OpenStudy (anonymous):

So, would I put the 27^3 as (3x)^3?

OpenStudy (anonymous):

Or do I leave it alone to put it in the rule?

OpenStudy (solomonzelman):

It seems like you didn't state the question correctly, what was your original question ?

OpenStudy (anonymous):

The original question is, factor the following polynomial completely. 27x^3+8

OpenStudy (solomonzelman):

then yes.... 27x^3 = (3x)^3 and thus 27x^3+8 is same as (3x)^3+2^3. Now use the rule.

OpenStudy (anonymous):

Oh! Thank you! I got the correct answer! :D

OpenStudy (solomonzelman):

Anytime